# A Systematic Literature Review of Spatio-Temporal Graph Neural Network Models for Time Series Forecasting and Classification

### v2

@@ 5.7.3 Types of models
Among the 43 papers in the “Generic” thematic group, 33 use a purely convolutional GNN approach, and four employ a purely attentional model. Five paper lack sufficient detail to be classified within the taxonomy chosen (Refs. [art_85, art_385, art_386, art_389, art_407]), and another one declares that it uses an aggregating approach that is simpler than either convolutional or attentional approaches [art_81]. Another paper proposes a so-called multivariate time series with dynamic graph neural ordinary differential equations (MTGODE) model, where the continuous dynamics of simplified graph propagation is described by an ordinary differential equation (ODE) [art_75]. This is not the only paper in this section using differential equations. In Ref. [art_401], two domain specific ODEs (one for traffic speed evolution, and the other for disease propagation) are integrated into the convolutional GNN model. The authors argue that this dynamics-informed model approach enhances robustness, allow easier generalization when the equation of the dynamics is known, and reduces the number of model parameters.Additionally, during inference, the model only requires a single data point, as the ODEs explicitly incorporate the immediate dynamics.
@@
As for the definition of the graph structure, the majority of these models use self-learning techniques. This is a straightforward choice, since all the papers in this group want to ensure the applicability of the model in different contexts, and the use of pre-defined rules for the definition of the graph limits the model’s universal applicability. More than half of the papers employ a dynamic graph structure, which adapts to the varying temporal patterns observed in the data.
@@
The two most commonly utilized loss functions are the MSE and the MAE. Since the graph structure is typically learned by the models themselves, many papers include a regularization term in the loss function for the optimization of the graph structure (e.g., Refs. [art_73, art_78, art_79, art_86]).
The majority of the papers specify the language and libraries used for the code, namely Python with PyTorch, PyTorch Geometric [PyTorch_Geometric], and Torch Spatiotemporal [Cini_Torch_Spatiotemporal_2022]. However, only slightly less than half of the papers provide a link to the source code. The links to the code repositories are listed in Tab. LABEL:t:link_codici_generic.
@@
Adaptive dependency learning neural network (ADLNN) [art_83] https://github.com/AbishekSriramulu/ADLGNN.git
Attention-adjusted Graph Spatio-Temporal Network (AGSTN) [art_464] https://github.com/l852888/AGSTN
Cross-attention temporal dynamic graph neural network (CATodyNet) [art_386] https://github.com/HQYWY/CATodyNet
Graph neural network with neural Granger causality (CauGNN) [art_77] https://github.com/RRRussell/CauGNN
Dynamic diffusion-variational GNN (DVGNN) [art_383] https://github.com/gorgen2020/DVGNN
Gformer [art_82] https://github.com/wxh453751461/Gformer
GraphSensor [art_395] https://github.com/jldx-gjq/Graphnet/tree/main/graphsensor-code-main
Multi-scale adaptive graph neural network (MAGNN) [art_80] https://github.com/shangzongjiang/MAGNN
Multi-task Dynamic Graph generation framework for multivariate time series forecasting (MDG) [art_397] https://github.com/gortwwh/MDG
Multivariate time series deep spatiotemporal forecasting model with a graph neural network (MDST-GNN) [art_90] https://github.com/yiminghzc/MDST-GNN
Multivariate Time Series Classification based on Fusion Features (MSTC_FF) [art_385] https://github.com/dumingsen/MTSC_FF
Multivariate time series with dynamic graph neural ordinary differential equations (MTGODE) [art_75] https://github.com/GRAND-Lab/MTGODE
Static and dynamic graph learning network (SDGL) [art_78] https://github.com/ZhuoLinLi-shu/SDGL
Sparse graph learning from spatiotemporal time series [art_72] https://github.com/andreacini/sparse-graph-learning
Temporal decomposition enhanced graph neural network for multivariate time series forecasting (TDG4MSF) [art_81] https://github.com/TYUT-Theta/MHZN.git
Temporal graph convolution and attention (T-GAN) [art_202] https://github.com/malei666666/T_GAN
Temporal Dynamic Graph Neural Network (TodyNet) [art_389] https://github.com/liuxz1011/TodyNet
Two GNN models with different graph structures [art_160] https://github.com/StefanBloemheuvel/graph_comparison
@@ 5.7.4 Benchmark models
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_73], [art_74], [art_78], [art_79], [art_80], [art_81], [art_83], [art_84], [art_88], [art_89], [art_384], [art_393], [art_397], [art_400], [art_405], [art_464]
Gaussian process (GP) [art_73], [art_74], [art_78], [art_80], [art_81], [art_83]
Historical average (HA) [art_78], [art_88], [art_400]
Vector autoregression (VAR) [art_73], [art_74], [art_77], [art_78], [art_79], [art_81], [art_84], [art_88], [art_384], [art_400], [art_405]
Non-GNN machine learning methods Autoformer [Autoformer] [art_74], [art_76], [art_85]
DeepAR [DeepAR] [art_69], [art_393]
DeepSleepNet [DeepSleepNet] [art_395], [art_403]
@@
Dual self-attention network for multivariate time series forecasting (DSANet) [DSANet] [art_73], [art_78], [art_405]
Feed-forward neural network (FNN) [art_84], [art_89]
Gated recurrent unit (GRU) [art_73], [art_78], [art_82], [art_88], [art_90], [art_384], [art_398], [art_404], [art_405]
Hybrid framework based on fully Dilated CNN (HyDCNN) [HyDCNN] [art_75], [art_83]
Informer [Informer] [art_74], [art_76], [art_82], [art_85], [art_397]
$k$-nearest neighbors (KNN) [art_84], [art_387]
LogTrans [LogTrans] [art_82], [art_85]
Long-short term memory network (LSTM) [art_71], [art_74], [art_79], [art_82], [art_84], [art_85], [art_86], [art_88], [art_89], [art_396], [art_397], [art_398], [art_400], [art_464]
Long-short term time series network (LSTNet) [LSTNet] [art_73], [art_74], [art_75], [art_77], [art_78], [art_80], [art_81], [art_82], [art_83], [art_85], [art_87], [art_89], [art_90], [art_387], [art_393]
Multi-Level Construal Neural Network (MLCNN) [MLCNN] [art_77], [art_87]
@@
Neural basis expansion analysis for interpretable time series forecasting (N-BEATS) [N-BEATS] [art_69], [art_85]
@@
Recurrent neural network with fully connected gated recurrent units (RNN-GRU) [art_73], [art_77], [art_78], [art_79], [art_80], [art_81], [art_83], [art_87], [art_405]
Reformer [Reformer] [art_76], [art_82], [art_85]
SCINet [SCINet] [art_74], [art_90]
Support vector machine (SVM) [art_84], [art_396], [art_464]
Temporal convolutional network (TCN) [TCN] [art_88], [art_387], [art_396], [art_397]
@@
Temporal pattern attention long-short term memory network (TPA-LSTM) [TPA-LSTM] [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_90], [art_405]
Transformer [attention] [art_71], [art_77], [art_393]
Hybrid of the multilayer perception and autoregressive model (VAR-MLP) [VAR-MLP] [art_73], [art_74], [art_78], [art_79], [art_80], [art_81], [art_83], [art_87], [art_90], [art_405]
GNN methods Adaptive graph convolutional recurrent network (AGCRN) [AGCRN] [art_73], [art_74], [art_78], [art_79], [art_80], [art_88], [art_384], [art_396], [art_400], [art_464]
@@
Attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN] [art_78], [art_79], [art_383], [art_384], [art_396], [art_397], [art_400], [art_464]
@@
Attention based Spatial-Temporal Graph Neural Network (ASTGNN) [ASTGNN] [art_74], [art_396]
@@
Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_73], [art_74], [art_75], [art_78], [art_79], [art_84], [art_86], [art_87], [art_88], [art_383], [art_384], [art_396], [art_397], [art_400], [art_405]
Dynamic Graph Convolutional Recurrent Network (DGCRN) [art_100] [art_397], [art_398]
@@
Dynamic and Multi-faceted SpatioTemporal Graph Convolution Network (DMSTGCN) [DMSTGCN] [art_396], [art_417]
@@
Dynamic Spatial-Temporal Aware Graph Neural Network (DSTAGNN) [DSTAGNN] [art_383], [art_384]
Graph auto encoder (GAE) [GAE] [art_70], [art_202]
Graph multi-attention network (GMAN) [GMAN] [art_73], [art_75], [art_86], [art_88]
Discrete graph structure learning for time series (GTS) [GTS] [art_72], [art_78], [art_384], [art_397], [art_401]
Graph WaveNet (GWN) [GWN] [art_74], [art_75], [art_79], [art_80], [art_86], [art_87], [art_396], [art_397], [art_417]
@@
Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_72], [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_86], [art_87], [art_88], [art_90], [art_393], [art_396], [art_397], [art_401], [art_405], [art_417]
@@
Multivariate Time series with dynamic Graph neural Ordinary Differential Equations (MTGODE) [art_75] [art_88], [art_401], [art_417]
Node2vec [node2vec] [art_70], [art_202]
@@
Spectral temporal graph neural network for multivariate time-series forecasting (StemGNN) [StemGNN] [art_397], [art_405]
Spatial-temporal fusion graph neural network (STFGNN) [STFGNN] [art_73], [art_86], [art_405]
Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_73], [art_75], [art_78], [art_79], [art_84], [art_86], [art_88], [art_384], [art_397], [art_400], [art_401]
Spatial-temporal graph ODE networks (STGODE) [STGODE] [art_75], [art_88], [art_383]
Spatio-temporal synchronous graph convolutional network (STSGCN) [STSGCN] [art_73], [art_74], [art_78], [art_387], [art_397], [art_400], [art_405]
Temporal graph convolutional network (T-GCN) [T-GCN] [art_383], [art_384], [art_393]
@@
Time zigzags at graph convolutional networks for time series forecasting (Z-GCNETs) [Z-GCNETs] [art_400], [art_405]
@@
This “Generic” thematic group includes a wide variety of benchmark models, also due to the large number of papers within this group. As for mathematical and statistical models, autoregressions and ARIMA models are very popular for their simplicity and interpretability. Among non-GNN machine learning models, the most widely used models are LSTM networks, the so-called temporal pattern attention long-short term memory network (TPA-LSTM) [TPA-LSTM], and the long-short term time series network (LSTNet) [LSTNet]. The TPA-LSTM model, introduced by Shih et al. in 2019, employs a set of filters to capture time-invariant temporal patterns, and an attention mechanism to identify relevant time series for multivariate forecasting. The source code for this benchmark model is available at https://github.com/gantheory/TPA-LSTM. The LSTNet model, proposed by Lai et al. in 2018, combines a convolution neural network and a recurrent neural network to capture short-term local dependencies among variables and identify long-term patterns. Additionally, it incorporates an autoregressive model that enhances the robustness of the deep learning approach to time series with significant scale fluctuations. The source code for the model is accessible at https://github.com/laiguokun/LSTNet. Among GNN models, the most common ones are DCRNN [DCRNN] and MTGNN [MTGNN] models, already discussed in the previous subsections.

### v4

@@ 5.7.3 Types of models
Among the 77 papers in the “Generic” thematic group, 56 use a purely convolutional GNN approach, and 10 employ a purely attentional model. Five paper lack sufficient detail to be classified within the taxonomy chosen (Refs. [186, 74, 113, 261, 140]), and another one declares that it uses an aggregating approach that is simpler than either convolutional or attentional approaches [287]. Moreover, in Ref. [59] the authors use an approach that can be proved to be equivalent to a convolutional model; Ref. [59] perform a general analysis that extends beyond a single model; Ref. [327] explores various models, all based on convolutional architectures; Ref. [300] proposes an architecture that can adopt either a convolutional or an attentional aggregation. Another paper proposes a so-called multivariate time series with dynamic graph neural ordinary differential equations (MTGODE) model, where the continuous dynamics of simplified graph propagation is described by an ordinary differential equation (ODE) [177]. This is not the only paper in this section using differential equations. In Ref. [360], two domain specific ODEs (one for traffic speed evolution, and the other for disease propagation) are integrated into the convolutional GNN model. The authors argue that this dynamics-informed model approach enhances robustness, allow easier generalization when the equation of the dynamics is known, and reduces the number of model parameters. Additionally, during inference, the model only requires a single data point, as the ODEs explicitly incorporate the immediate dynamics.
@@
As for the definition of the graph structure, the majority of these models use self-learning techniques. This is a straightforward choice, since all the papers in this group want to ensure the applicability of the model in different contexts, and the use of pre-defined rules for the definition of the graph limits the model’s universal applicability and negatively impacts knowledge transfer. More than half of the papers employ a dynamic graph structure, which adapts to the varying temporal patterns observed in the data.
In Ref. [451] the authors introduce a novel data structure called hypervariate graph, where each time serie values is represented as a graph node, and slicing windows are modeled as fully connected space-time graphs. Unlike traditional approaches that treat spatial and temporal aspects separately, according to the authors, this approach captures spatiotemporal correlations in an integrated manner.
@@
The two most commonly utilized loss functions are the MSE and the MAE. Since the graph structure is typically learned by the models themselves, many papers include a regularization term in the loss function for the optimization of the graph structure (e.g., Refs. [241, 243, 118, 52, 338]).
The majority of the papers specify the language and libraries used for the code, namely Python with PyTorch, PyTorch Geometric [91], and Torch Spatiotemporal [58]. However, only slightly more than half of the papers provide a link to the source code. The links to the code repositories are listed in Tab. LABEL:t:link_codici_generic. Ref. [59] is a study that does not introduce a new model but rather investigates the relationship between global and local patterns in graph-based spatiotemporal forecasting. The authors provide access to their source code at https://github.com/Graph-Machine-Learning-Group/taming-local-effects-stgnns. Additionally, this group includes the work that introduced the PyTorch Geometric Temporal library [327], a framework for dynamic graph learning. The corresponding source code is publicly available at https://github.com/benedekrozemberczki/pytorch_geometric_temporal. In addition, the authors compare various open-source libraries for geometric deep learning.
@@
Adaptive dependency learning neural network (ADLNN) [355] https://github.com/AbishekSriramulu/ADLGNN.git
Attention-adjusted graph spatio-temporal network (AGSTN) [277] https://github.com/l852888/AGSTN
Cross-attention temporal dynamic graph neural network (CATodyNet) [113] https://github.com/HQYWY/CATodyNet
Graph neural network with neural Granger causality (CauGNN) [76] https://github.com/RRRussell/CauGNN
Spatio-temporal contrastive learning (CL4ST) [368] https://github.com/HKUDS/CL4ST
CrossGNN [153] https://github.com/hqh0728/CrossGNN
Dynamic diffusion-variational GNN (DVGNN) [244] https://github.com/gorgen2020/DVGNN
Evolving multi-scale graph neural network (ESG) [448] https://github.com/LiuZH-19/ESG
Fourier graph neural network (FourierGNN) [451] https://github.com/aikunyi/FourierGNN
Graph interpolation attention recursive network (GinAR) [458] https://github.com/ChengqingYu/GinAR
Gformer [395] https://github.com/wxh453751461/Gformer
Graphs for forecasting irregularly sampled time series with missing values (GraFITi) [436] https://github.com/yalavarthivk/GraFITi
GraphSensor [100] https://github.com/jldx-gjq/Graphnet/tree/main/graphsensor-code-main
Graph for time series (GTS) [338] https://github.com/chaoshangcs/GTS
HAR-Sensor [289] https://github.com/riktimmondal/HAR-Sensor
Hierarchical downsampling time-then-space (HD-TTS) [286] https://github.com/marshka/hdtts
HyperTime [456] https://gitfront.io/r/user-9180183/WXPgc7aQjSKE/HyperTime/
Multi-scale adaptive graph neural network (MAGNN) [40] https://github.com/shangzongjiang/MAGNN
Multi-task dynamic graph generation framework for multivariate time series forecasting (MDG) [152] https://github.com/gortwwh/MDG
Multivariate time series deep spatiotemporal forecasting model with a graph neural network (MDST-GNN) [138] https://github.com/yiminghzc/MDST-GNN
Multivariate time series classification based on fusion features (MSTC_FF) [74] https://github.com/dumingsen/MTSC_FF
Multivariate time series forecasting with graph neural networks (MTGNN) [424] https://github.com/nnzhan/MTGNN
Multivariate time series with dynamic graph neural ordinary differential equations (MTGODE) [177] https://github.com/GRAND-Lab/MTGODE
RAINDROP [499] https://github.com/mims-harvard/Raindrop
Static and dynamic graph learning network (SDGL) [243] https://github.com/ZhuoLinLi-shu/SDGL
Similarity-aware time-series classification (SimTSC) [472] https://github.com/daochenzha/SimTSC
Sparse graph learning from spatiotemporal time series [60] https://github.com/andreacini/sparse-graph-learning
Spatial relation decomposition (SRD) [83] https://github.com/Arthur-Null/SRD/tree/main
Spatio-temporal graph neural networks (STExplainer) [369] https://github.com/HKUDS/STExplainer
Time-aware multipersistence spatio-supra graph convolutional network (TAMP-S2GCNet) [51] https://www.dropbox.com/sh/n0ajd5l0tdeyb80/AABGn-ejfV1YtRwjf_L0AOsNa?dl=0
Temporal decomposition enhanced graph neural network for multivariate time series forecasting (TDG4MSF) [287] https://github.com/TYUT-Theta/MHZN.git
Temporal graph convolution and attention (T-GAN) [155] https://github.com/malei666666/T_GAN
Temporal graph neural networks for irregular data (TGNN4I) [300] https://github.com/joeloskarsson/tgnn4i
Temporal dynamic graph neural network (TodyNet) [261] https://github.com/liuxz1011/TodyNet
Transformable patching graph neural network (T-PATCHGNN) [491] https://github.com/usail-hkust/t-PatchGNN
Temporal polynomial graph neural network (TPGNN) [271] https://github.com/zyplanet/TPGNN
Two GNN models with different graph structures [15] https://github.com/StefanBloemheuvel/graph_comparison
Unified Spatio-Temporal Diffusion Models (USTD) [143] https://github.com/hjf1997/USTD
Weight-bounded-GAT [505] https://github.com/zzh237/Weight-bounded-GAT
Zigzag filtration curve based supra-hodge convolution networks (ZFC-SHCN) [48] https://github.com/zfcshcn/ZFC-SHCN
Time zigzags at graph convolutional network (Z-GCNET) [50] https://github.com/Z-GCNETs/Z-GCNETs
@@ 5.7.4 Benchmark models
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [241], [99], [243], [118], [40], [287], [355], [47], [473], [57], [231], [152], [390], [240], [450], [338], [83], [448], [277], [424], [61], [271], [49]
Gaussian process (GP) [241], [99], [243], [40], [287], [355], [83], [448], [424], [61], [271]
Historical average (HA) [243], [473], [390], [338], [279], [61], [50], [369], [368]
Vector autoregression (VAR) [241], [99], [76], [243], [118], [287], [47], [473], [390], [240], [432], [338], [50], [451], [356], [369], [368]
Non-GNN machine learning methods Autoformer [414] [99], [504], [186], [432], [83], [153], [451]
Crossformer [503] [99], [491], [153]
DeepAR [329] [36], [231], [409]
Attentive hierarchical recurrent networks for crime prediction (DeepCrime) [149] [369], [368]
DeepGLO [335] [432], [51], [26], [48], [451], [49]
DeepSleepNet [364] [100], [401]
DeepState [321] [51], [26], [48], [49]
DLinear [469] [504], [436], [491], [153]
Deep multi-view spatial-temporal network (DMVST-Net) [447] [450], [49]
@@
Dual self-attention network for multivariate time series forecasting (DSANet) [157] [241], [243], [240], [50], [368]
Exponential smoothing transformers for time-series forecasting (ETSformer) [412] [504], [153]
Frequency enhanced decomposed transformer (FEDformer) [522] [99], [432], [436], [153]
Feed-forward neural network (FNN) [47], [57], [450], [338], [472]
Gated recurrent unit (GRU) [241], [243], [395], [473], [138], [453], [337], [240], [286], [499], [491], [50], [436], [356], [300]
Hybrid framework based on fully Dilated CNN (HyDCNN) [233] [177], [355]
InceptionTime [162] [83], [472]
Informer [520] [99], [504], [395], [186], [152], [432], [436], [279], [61], [271], [451], [356]
$k$-nearest neighbors (KNN) [47], [396], [456], [472]
LogTrans [294] [395], [186]
Long-short term memory network (LSTM) [129], [99], [118], [395], [47], [186], [52], [473], [57], [137], [152], [453], [390], [338], [436], [279], [277], [50], [51], [26], [271], [48], [356], [49]
Long-short term time series network (LSTNet) [202] [241], [99], [177], [76], [243], [40], [287], [395], [355], [186], [397], [57], [138], [396], [231], [432], [83], [448], [424], [61], [51], [26], [271], [48], [451], [356], [49], [268]
Multi-level construal neural network (MLCNN) [54] [76], [397]
Multi-time attention networks (mTANs) [349] [436], [499], [491]
@@
Neural basis expansion analysis for interpretable time series forecasting (N-BEATS) [299] [36], [186], [51], [26], [48], [49]
Channel-independent patch time series Transformer (PatchTST) [295] [491], [153]
Reformer [191] [504], [395], [186], [432], [451]
Recurrent neural network (RNN) [57], [59]
@@
Recurrent neural network with fully connected gated recurrent units (RNN-GRU) [241], [76], [243], [118], [40], [287], [355], [397], [240], [448], [424], [271]
Rocket [64] [140], [83]
SCINet [264] [99], [138]
Set functions for time series (SeFT) [141] [499], [491]
State frequency memory (SFM) [484] [432], [51], [26], [48], [451]
Shapelet-neural network (ShapeNet) [215] [113], [261], [456]
Spatial-temporal dynamic network (STDN) [446] [450], [369], [49], [368]
Spatial-temporal identity (STID) [340] [99], [458]
Spatial-temporal transformer network (STtrans) [418] [369], [368]
Spatio-temporal residual networks (ST-ResNet) [480] [369], [49], [368]
Support vector machine (SVM) [47], [137], [338], [277], [49], [368]
TapNet [495] [74], [261], [456], [472]
Temporal convolutional network (TCN) [9] [473], [396], [137], [152], [432], [51], [26], [48], [451], [49]
Temporal 2D-variation modeling for general time series analysis (TimesNet) [413] [504], [491], [456], [153]
@@
Temporal pattern attention long-short term memory network (TPA-LSTM) [347] [241], [99], [177], [243], [118], [40], [287], [355], [138], [240], [448], [424], [61], [271], [268]
TimeGrad [322] [143], [409]
Transformer [375] [129], [76], [231], [499], [300]
Hybrid of the multilayer perception and autoregressive model (VAR-MLP) [485] [241], [99], [243], [118], [40], [287], [355], [397], [138], [240], [448], [424], [61], [271], [268]
WEASEL + multivariate unsupervised symbols and derivatives (WEASEL+MUSE) [333] [74], [113], [261], [456]
Extreme gradient boosting (XGBoost) [450], [448], [279]
GNN methods Adaptive graph convolutional recurrent network (AGCRN) [8] [241], [99], [243], [118], [40], [473], [137], [390], [432], [286], [279], [50], [51], [48], [451], [59], [369], [49], [327], [368]
@@
Attention-based spatial-temporal graph convolutional network (ASTGCN) [116] [243], [118], [244], [137], [152], [390], [277], [50], [369], [505], [368]
@@
Attention based spatial-temporal graph neural network (ASTGNN) [117] [99], [137]
@@
Coupled layer-wise convolutional recurrent neural network (CCRNN) [449] [448], [279]
CrossGNN [153] [491], [153]
Diffusion convolutional recurrent neural network (DCRNN) [237] [241], [99], [177], [243], [118], [47], [52], [397], [473], [244], [137], [152], [390], [240], [338], [286], [83], [448], [279], [277], [424], [50], [51], [271], [48], [59], [143], [409], [369], [49], [327], [368]
Dynamic graph convolutional recurrent network (DGCRN) [212] [152], [453]
@@
Dynamic and multi-faceted spatiotemporal graph convolution network (DMSTGCN) [127] [137], [443], [83], [369], [368]
@@
Dynamic spatial-temporal aware graph neural network (DSTAGNN) [203] [244], [369]
Evolving multi-scale graph neural network (ESG) [448] [99], [279]
Graph auto encoder (GAE) [189] [232], [155]
Graph deviation network (GDN) [66] [396], [505]
Graph multi-attention network (GMAN) [513] [241], [177], [52], [473], [424], [369], [505], [368]
Graph-based multi-step dependency relation (GMSDR) [259] [143], [409], [369], [368]
Discrete graph structure learning for time series (GTS) [338] [60], [243], [152], [360], [448], [279]
Graph WaveNet (GWN) [425] [99], [177], [118], [40], [52], [397], [137], [152], [443], [432], [286], [491], [279], [424], [50], [51], [26], [271], [48], [451], [59], [143], [369], [49], [368]
Hypergraph convolutional recurrent neural network (HGC-RNN) [450] [390], [49]
Long short-term graph convolutional networks (LSGCN) [154] [50], [368]
Multi-range attentive bicomponent graph convolutional network (MRA-BGCN) [44] [52], [424]
@@
Multivariate time series forecasting with graph neural network (MTGNN) [424] [60], [241], [99], [177], [243], [118], [40], [287], [355], [52], [397], [473], [138], [231], [137], [152], [360], [240], [443], [432], [491], [83], [448], [61], [271], [153], [451], [268], [505]
@@
Multivariate time series with dynamic graph neural ordinary differential equations (MTGODE) [177] [473], [360], [443]
Node2vec [110] [232], [155]
@@
Spectral temporal graph neural network for multivariate time-series forecasting (StemGNN) [26] [152], [240], [432], [491], [448], [51], [26], [271], [48], [451], [369], [49], [368]
Spatial-temporal fusion graph neural network (STFGNN) [228] [241], [52], [240], [369], [368]
Spatio-temporal graph convolutional network (STGCN) [457] [241], [177], [243], [118], [47], [52], [473], [152], [390], [360], [448], [424], [50], [51], [26], [271], [48], [143], [409], [369], [49], [505], [368]
Spatio-temporal graph neural controlled differential equation (STG-NCDE) [56] [177], [143], [409], [369], [368]
Spatial-temporal graph ODE networks (STGODE) [84] [177], [473], [244], [143], [369], [368]
Deep-meta-learning based model (ST-MetaNet) [306] [52], [424], [369], [49], [368]
Spatiotemporal multi-graph convolution network (ST-MGCN) [101] [277], [49]
Spatio-temporal synchronous graph convolutional network (STSGCN) [353] [241], [99], [243], [396], [152], [390], [240], [448], [50], [143], [369], [368]
Spatial-temporal sequential hypergraph network (ST-SHN) [427] [369], [368]
Time-aware multipersistence spatio-supra graph convolutional network (TAMP-S2GCNet) [51] [369], [368]
Temporal graph convolutional network (T-GCN) [508] [244], [231], [327]
Temporal dynamic graph neural network (TodyNet) [261] [113], [140], [456]
@@
Time zigzags at graph convolutional networks for time series forecasting (Z-GCNETs) [50] [390], [240], [51], [26], [48], [369], [49], [368]
@@
This “Generic” thematic group includes a wide variety of benchmark models, also due to the large number of papers within this group. As for mathematical and statistical models, autoregressions and ARIMA models are very popular for their simplicity and interpretability. Among non-GNN machine learning models, the most widely used models are LSTM networks, the so-called temporal pattern attention long-short term memory network (TPA-LSTM) [347], and the long-short term time series network (LSTNet) [202]. The TPA-LSTM model, introduced by Shih et al. in 2019, employs a set of filters to capture time-invariant temporal patterns, and an attention mechanism to identify relevant time series for multivariate forecasting. The source code for this benchmark model is available at https://github.com/gantheory/TPA-LSTM. The LSTNet model, proposed by Lai et al. in 2018, combines a convolution neural network and a recurrent neural network to capture short-term local dependencies among variables and identify long-term patterns. Additionally, it incorporates an autoregressive model that enhances the robustness of the deep learning approach to time series with significant scale fluctuations. The source code for the model is accessible at https://github.com/laiguokun/LSTNet. Among GNN models, the most common ones are DCRNN [237], MTGNN [424], and GWN [425] models, already discussed in the previous subsections.